Excerpts from Dowling Company's December 31, 2021 and 2020, financial statements and key ratios are presented below (all numbers are in millions): 2021 2020 Accounts receivable (net) $ 21 $ 13 Net sales $ 125 $ 110 Cost of goods sold $ 70 $ 65 Net income $ 21 $ 16 Inventory turnover 5.70 Return on assets 11.50 % Equity multiplier 2.46 Dowling's 2021 average collection period is: (Round your answer to 2 decimal places.)
Answer #1
Average accounts receivable = ( Beginning account receivable + Ending accounts receivable ) / 2 = ( 13000000 + 21000000 ) / 2 17000000
Average collection period = ( Average accounts receivable / Net sales ) * Days in a year = ( 17000000 / 125000000 ) * 365 49.64 days